If 54.5 mol of an ideal gas occupies 31.5 L at 17.00 ∘C, what is the pressure of the gas?

?=   atm
Answer #1

given) P=D (we have to find) T=17°C = (19+273) K =5290k ~= 31.5L n = 54.5 mol - Now we Pr=nRT We P= (MRT) R = 0.08206 L-atm/m